# Physically-Consistent Parameter Identification of Robots in Contact

### v1

@@ Abstract
Accurate inertial parameter identification is crucial for the simulation and control of robots encountering intermittent contact with the environment. Classically, robots’ inertial parameters are obtained from CAD models that are not precise (and sometimes not available, e.g., Spot from Boston Dynamics), hence requiring identification. To do that, existing methods require access to contact force measurement, a modality not present in modern quadruped and humanoid robots. This paper presents an alternative technique that utilizes joint current/torque measurements —a standard sensing modality in modern robots— to identify inertial parameters without requiring direct contact force measurements. By projecting the whole-body dynamics into the null space of contact constraints, we eliminate the dependency on contact forces and reformulate the identification problem as a linear matrix inequality that can handle physical and geometrical constraints. We compare our proposed method against a common black-box identification mrethod using a deep neural network and show that incorporating physical consistency significantly improves the sample efficiency and generalizability of the model. Finally, we validate our method on the Spot quadruped robot across various locomotion tasks, showcasing its accuracy and generalizability in real-world scenarios over different gaits.
@@ I Introduction
Classically, system identification of legged robots have largely relied on full Force/Torque (F/T) sensing, utilizing either onboard F/T sensors at the end-effectors [5] or data collected from F/T plates during locomotion [6]. However, these methods have practical limitations; not all legged robots are endowed with full F/T sensors, and F/T plates are often impractical for real-world applications. Indeed, most state-of-the-art quadrupedal [7, 8, 9] and bipedal [10, 11, 12] platforms can only measure or estimate joint torques (from motor current), without F/T sensors at end-effectors. Hence, a new inertial parameters identification framework that relies only on joint torque data is desired.
Figure 1: Whole-body inertial parameter identification of floating-base robots using constraint-consistent dynamics with joint torque measurements and constrained convex optimization.
@@ II Related Work
With the development of more complex robotic systems, particularly legged robots with floating bases, research has shifted toward addressing the unique challenges posed by these platforms, such as contact switches, limited range of motion during identification, and under-actuation of the robot base. The identifiability of a minimal set of inertial parameters, also known as base inertial parameters, was demonstrated in [6] by leveraging only the unactuated dynamics of the floating base. Later works, like [5], used the whole-body dynamics of the floating base for parameter identification, assuming the availability of full contact F/T sensing and joint torque measurements. In addressing the challenge of inertial parameter estimation with partial F/T sensing, [17] identified a subset of sensors and utilized singular value decomposition (SVD) techniques to identify this subset for identifiability of base parameters. Contrary to [5], [6] that require F/T sensor for parameter identification, our approach relies solely on joint torque measurements, eliminating the need for contact force data. Furthermore, unlike our formulation, these techniques solve an unconstrained linear least squares problem, hence they often struggle with ensuring the physical feasibility of the identified parameters.
@@
More recently, researchers have recognized that inertial parameters reside on a curved Riemannian space [21], which complicates the optimization process. Regularization techniques based on coordinate-invariant distances [22] have been shown to improve accuracy and robustness in the identification process, however leading to non-convex and computationally expensive optimization problems. To overcome these computational challenges while maintaining robustness against noise, [23] proposed a second-order approximation of the geodesic distance as a regularization term, achieving convexity without sacrificing performance. We leverage this regularization term to improve the robustness of our identification while being computationally efficient.
@@ III-C Motion and Constraint Dynamics
IV Inertial Parameters Identification

### v2

@@ Abstract
Accurate inertial parameter identification is crucial for the simulation and control of robots encountering intermittent contacts with the environment. Classically, robots’ inertial parameters are obtained from CAD models that are not precise (and sometimes not available, e.g., Spot from Boston Dynamics), hence requiring identification. To do that, existing methods require access to contact force measurement, a modality not present in modern quadruped and humanoid robots. This paper presents an alternative technique that utilizes joint current/torque measurements —a standard sensing modality in modern robots— to identify inertial parameters without requiring direct contact force measurements. By projecting the whole-body dynamics into the null space of contact constraints, we eliminate the dependency on contact forces and reformulate the identification problem as a linear matrix inequality that can handle physical and geometrical constraints. We compare our proposed method against a common black-box identification method using a deep neural network and show that incorporating physical consistency significantly improves the sample efficiency and generalizability of the model. Finally, we validate our method on the Spot quadruped robot across various locomotion tasks, showcasing its accuracy and generalizability in real-world scenarios over different gaits.
@@ I Introduction
Classically, system identification of legged robots have largely relied on full Force/Torque (F/T) sensing, utilizing either onboard F/T sensors at the end-effectors [5] or data collected from F/T plates during locomotion [6]. However, these methods have practical limitations; not all legged robots are endowed with full F/T sensors, and F/T plates are often impractical for real-world applications. Indeed, most state-of-the-art quadrupedal [7, 8, 9] and bipedal [10, 11, 12] platforms can only measure or estimate joint torques (from motor current), without F/T sensors at end-effectors. Hence, a new inertial parameter identification framework that relies only on joint torque data is desired.
Figure 1: Whole-body inertial parameter identification —mass, center of mass and moment of inertia matrix— of floating-base robots using constraint-consistent dynamics with joint torque measurements and constrained convex optimization.
@@ II Related Work
With the development of more complex robotic systems, particularly legged robots with floating bases, research has shifted toward addressing the unique challenges posed by these platforms, such as contact switches, limited range of motion during identification, and under-actuation of the robot base. The identifiability of a minimal set of inertial parameters, also known as base inertial parameters, was demonstrated in [6] by leveraging only the unactuated dynamics of the floating base. Later works, like [5], used the whole-body dynamics of the floating base for parameter identification, assuming the availability of full contact F/T sensing and joint torque measurements. In addressing the challenge of inertial parameter estimation with partial F/T sensing, [17] identified a subset of sensors and utilized Singular Value Decomposition (SVD) techniques to identify this subset for identifiability of base parameters. Unlike [5] and [6], which require F/T sensors for parameter identification, our approach relies solely on joint torque measurements, eliminating the need for contact force data. Although [18] also employs the same approach to remove the dependency on contact forces, their work focuses on learning residual model errors for improved inverse dynamics control rather than identifying inertial parameters. Moreover, by formulating the identification problem as a constrained optimization using LMI formalism, we ensure physical consistency—an aspect often neglected in traditional unconstrained least squares methods.
@@
More recently, researchers have recognized that inertial parameters reside on a curved Riemannian manifold [22], which complicates the optimization process. In their work, the authors have shown that regularization techniques based on coordinate-invariant distances improves accuracy and robustness in the identification process, however leading to non-convex and computationally expensive optimization problems. To overcome these computational challenges while maintaining robustness against noise, [23] proposed a second-order approximation of the geodesic distance as a regularization term, achieving convexity without sacrificing performance. We leverage this regularization term to improve the robustness of our identification while being computationally efficient.
@@ III-C Motion and Constraint Dynamics
IV Inertial Parameter Identification